OpenStudy (anonymous):

integrate from 0 to 1.5 [x^2]dx where [x] is the greatest integer function

OpenStudy (watchmath):

\(\int_0^{1.5}[x^2]\,dx=\int_0^10\,dx+\int_1^{\sqrt{2}} 1\,dx+\int_{\sqrt{2}}^{1.5} 2\,dx\)